Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-8x – 4y = -68
x + 3y = 16

Answer by ewatrrr(24785)

 
Hi,
-8x – 4y = -68
x + 3y = 16

-8x – 4y = -68
8x + 24y = 128 |Multiplying 2nd EQ thru by 8 to eliminate x
20y = 60
y = 3 and x = 7 (x = -3*3 + 16)
Ordered pair Pt(7,3)is the solution for this INDEPENDENT system:See below
Lines intersect at Pt(7,3)

A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution. ***
A dependent system has infinitely many solutions.